OpenStudy (anonymous):

solve 3/4 x +18 = 18

OpenStudy (anonymous):

3/4x + 18 = 18 -18 -18 3/4x = 0 x= 0

OpenStudy (anonymous):

is this a no solution problem?

sammixboo (sammixboo):

\(\sf \large \color{magenta}{3/4x~+~18~=~18}\) \(\sf \large \color{magenta}{3/4x~+~18~-~18=~18~-~18}\) \(\sf \large \color{magenta}{3/4x~=~0}\) \(\sf \large \color{magenta}{x~=~0}\)

OpenStudy (anonymous):

your solution is 0. 0 is a perfectly fine number. An no solution answer would be x/0 which doesn't exist.